# Entanglement behavior and localization properties in monitored fermion systems

### v1

@@ IV.1 Tight-binding chain with onsite dephasing
where $J$ denotes the hopping strength and $\hat{n}_{j}=\hat{c}^{\dagger}_{j}\hat{c}_{j}$ is the onsite fermion number operator. Here and in the other considered one-dimensional models, we adopt periodic boundary conditions by assuming $\hat{c}^{(\dagger)}_{L+1}\equiv\hat{c}^{(\dagger)}_{1}$. With reference to Eq. (7), we have $p_{j}=0$ and $q_{j}=1$. This system possesses a $U(1)$ symmetry, corresponding to the conservation of the total number of fermions, $\hat{N}=\sum_{j}\hat{n}_{j}$. In this case, the unraveled state |$\ket{\psi(t)}$ can be always cast in a Slater determinant form [44, 65]
|$\ket{\psi(t)}=\prod_{k=1}^{N}\bigg{[}\,\sum_{j=1}^{L}\big{[}U_{t}\big{]}_{jk}\,\hat{c}_{j}^{\dagger}\bigg{]}\ket{\Omega}\,,$ (16)
@@
For this model, the authors of Ref. [44] showed the existence of an area-law phase for the asymptotic EE. More recently, the existence of a transition from area- to logarithm-law has been first claimed [52, 88] and then challenged. In fact, through the replica trick within a Keldysh path-integral formalism, it has been suggested that only the area-law phase exists, while the logarithm-law phase should be just a finite-size crossover, due to the exponential growth of a localization length with the inverse measurement strength [49].
Figure 1: The asymptotic averaged EE for the model in Eqs. (15). (a) Some examples with the behavior of $\overline{S}_{L/2}$ versus $L$ (circles) for different values of $\gamma$, together with the corresponding fits with Eq. (1) (continuous lines). (b) The fit parameter $b$ versus $\gamma$ (here and in the next figures, dashed lines are just to guide the eye). We simulate the time evolution until $t_{f}=6\times 10^{6}$, with a step $\delta t=0.01$. Here and in the next figures, we will always work in units of $J=1$.
Figure 1(a) displays numerical results for $\overline{S}_{L/2}$ versus $L$ (circles), for some values of the measurement strength $\gamma$, and the corresponding fit of Eq. (1) (continuous lines). We observe an accurate agreement between the two. Analyzing in more detail the fitting parameter $b$ versus $\gamma$ [Fig. 1(b)], we can see that, for $\gamma\gtrsim 0.4$, $b$ becomes larger than one. As expected, in correspondence of these measurement rates, the EE displays an area-law scaling. On the other hand, for $\gamma<0.4$, our results predict a power-law increase with an exponent slightly smaller than one, a behavior that might be mistaken for a logarithm, as found in Refs. [52, 88]. It is worth noticing that the value of $b$ reported in Fig. 1(b) is always very close to $1$ and exhibits a nonmonotonic behavior in $\gamma$, which is not observed in the other models. For such reason, in this case is very difficult to discriminate whether this $b<1$ regime corresponds to a genuine logarithm law or is just a finite-size effect, thus heralding an area-law behavior at larger system sizes.
@@ IV.2 Kitaev chain with onsite dephasing
|$\ket{\psi(t)}={\mathcal{N}}_{t}\,\exp{\bigg{[}\frac{1}{2}\sum_{j_{1},j_{2}=1}^{L}\big{[}{Z}_{t}\big{]}_{j_{1}j_{2}}{\hat{c}^{\dagger}}_{j_{1}}{\hat{c}^{\dagger}}_{j_{2}}\bigg{]}}\,|0\rangle\,,$ (18)
where ${\cal N}_{t}$ is a normalization prefactor and ${Z}_{t}$ is an antisymmetric $L\times L$ matrix that can be written as ${Z}_{t}=[{U}_{t}^{\dagger}]^{-1}\,{V}_{t}^{\dagger}$. The $U_{t}$ and $V_{t}$ can be cast as the submatrices of a Bogolibuov rotation allowing to construct the fermionic operators that annihilate the unraveled state (18) and obey linear differential equations [90, 75]. The interpretation as a Bogoliubov rotation is valid if ${U}_{t}$ and ${V}_{t}$ obey a unitarity condition, a constraint that can be restored (keeping ${Z}_{t}$ unchanged) by using a QR decomposition [75]. One can therefore restrict to study the dynamics of the two $L\times L$ matrices ${U}_{t}$ and ${V}_{t}$, keeping a polynomial scaling of the problem complexity and thus allowing the numerics to reach systems with a few hundreds of sites. Previous works [53, 56, 57] have shown that the asymptotic EE exhibits a transition from an area-law scaling to a logarithmic one: The transition point depends both on the measurement strength and on the parameter $h$ in the Hamiltonian (17a).
@@
Figure 2: The EE for the model in Eqs. (17). (a) $\overline{S}_{L/4}$ versus $L$ (circles) for different values of $h$ and fixed $\gamma=1.5$, together with the corresponding fits with Eq. (1) (continuous lines). (b) The fit parameter $b$ versus $h$. We evolve up to $t_{f}=60$ with a step $\delta t=0.05$.
@@ IV.3 Kitaev chain with long-range dissipators
Also in this model the $\mathbb{Z}_{2}$ symmetry associated to the parity is preserved and, due to the particular structure of the measurement operators, the quantum-state-diffusion dynamics preserves the Gaussianity of the unraveled state |$\ket{\psi(t)}$, that can be cast in the form Eq. (18). Previous numerical investigations of the dynamics of this model already showed the emergence of three parameter regions where the EE behaves distinctly, ranging from volume-law, to area-law, as well as to intermediate subvolume-law scaling with the system size [75].
All these regimes can be recognized by fitting the asymptotic averaged EE with Eq. (1). To show this fact, we concentrate on the case $\gamma=0.1$ and $h=0.5$. In Fig. 3(a) we show numerical data for different power-law exponents $\alpha$ (circles) and the corresponding fit (lines), that nicely reproduces all the curves. In Fig. 3(b) we show the parameter $b$ vs $\alpha$. The shaded areas locate the two crossover regions, respectively at $0.5\lesssim\alpha^{\star}_{1}\lesssim 1$ and at $\alpha^{\star}_{2}\approx 3.2$, discussed in Ref. [75]. We expect a volume-law behavior for $\alpha<\alpha^{\star}_{1}$ and, in fact, we find $b\approx 0$. In contrast, for $\alpha>\alpha^{\star}_{2}$, we expect an area-law regime and, in fact, we get $b>1$. Finally, in the intermediate region we expect a subvolume growth in $L$, confirmed by an exponent $b\approx 0.8$ which is roughly constant in all the region. Regarding the last observation, we point out that in Ref. [75] the observed subvolume growth was faster than the logarithmic one. Moreover, no analytical function to describe the behavior of the entanglement with the system size was proposed there, while the fitting function Eq. (1) we suggest here well describes this behavior. Even for this model, qualitatively analogous considerations apply for other values of $\gamma$.
Figure 3: The EE for the model in Eqs. (17a) and (LABEL:Z2_int_lr:eqn). (a) $\overline{S}_{L/2}$ versus $L$ (circles) for different values of $\alpha$ and fixed $\gamma=0.1$, $h=0.5$, together with the corresponding fits with Eq. (1) (continuous lines). (b) The fit parameter $b$ versus $\alpha$. We simulate the time evolution until $t_{f}=3\times 10^{3}$, with a step $\delta t=0.005$. The shaded areas locate the two crossover regions found in Ref. [75] for the same set of parameters used here.
@@ V Nonintegrable models
In Sec. V.1 we describe how $\overline{S}_{\ell}$ versus $\gamma$ can be fitted by the function in Eq. (20) for the $t$-$V$ model with onsite dephasing, while in Sec. V.2 we do the same for the SYK model. In Sec. V.3, we discuss how the parameters $K,\,Q,\,\beta$ depend on the size $L$. This analysis shows that, in both cases, the dependence of $\overline{S}_{\ell}$ on $L$ is of the same form as in Eq. (1); this finding provides us with the rationale for fitting $\overline{S}_{\ell}$ versus $\gamma$ with a non intuitive function as Eq. (20). As a last stage to understand nonintegrable models, in Sec. V.4 we focus on their localization properties, by checking for the scaling of the IPR with the dimension of the Hilbert space. We predict an anomalous delocalization behavior which is apparently not related to the EE.
@@ V.1 Staggered $t$-$V$ model with onsite dephasing
Figure 4: The EE for the model in Eqs. (21). Some examples of $\overline{S}_{L/2}$ versus $\gamma$ (circles), for various system sizes up to $L=20$ (see legend), together with the corresponding fits with the generalized Lorentzian function in Eq. (20) (continuous lines). We simulate the time evolution until $t_{f}=2\times 10^{3}$ with a step $\delta t=0.01$, while we set $W=V=1$.
Figure 4 display our numerical results for the asymptotic averaged EE $\overline{S}_{L/2}$ versus the measurement strength $\gamma$ (circles) and the corresponding fit obtained with Eq. (20) (continuous lines). We can see that the latter performs well over a range of $\gamma\in[8\times 10^{-3},4]$ corresponding to more than two orders of magnitude.
@@ V.2 SYK model with onsite dephasing
The SYK Hamiltonian is a fermionic long-range interacting lattice model, being characterized by random four-particle interactions. [93, 94] Adding dissipation in the form of local dephasing, as in Eq. (15), the model can be written as
@@
The SYK model is elusive to perturbative treatments at any energy scale, lying far outside the quasiparticle paradigm. In fact, it is known to be a paradigm for quantum chaos, displaying fast scrambling [98, 99], a nonzero entropy density at vanishing temperature [96], and exhibiting a volume-law bipartite EE for all the eigenstates (even for the ground state) [100, 101]. In the context of entanglement transitions, a related model of Brownian SYK chains subject to continuous monitoring has been considered in Ref. [83].
Figure 5: The EE for the model in Eqs. (22). We show some examples of $\overline{S}_{L/2}$ versus $\gamma$ (circles), for various sizes up to $L=20$ (see legend), together with the corresponding fits with Eq. (20) (continuous lines). We simulate the time evolution until $t_{f}=3.4\times 10^{2}$, with a step $\delta t=0.01$, and fix $J=1$.
In Fig. 5 we show the numerical results for the averaged asymptotic EE versus the measurement strength (circles) and the fit obtained with Eq. (20) (continuous lines), displaying a good agreement between the two over the same range of $\gamma\in[8\times 10^{-3},4]$, also for this model.
@@ V.3 Discussion
The results showed in the previous subsection cannot help too much in determining the asymptotic properties of the EE. However, some information can be deduced by looking at the other fitting parameters. In Fig. 6, we show the behaviors of $\beta$ vs $L$ (a), $K$ vs $L$ (b), and of $\ln Q$ vs $\ln L$ (c), for both the $t$-$V$ (orange) and the SYK (green) models.
@@
we find $x=1.023\pm 0.008$ for the $t$-$V$ model, while $x=0.955\pm 0.039$ for the SYK model. This result is not surprising, as $K$ is the value of the EE in the $\gamma\to 0$ limit, corresponding to the absence of an environment. This is known from literature to exhibit a volume law (see, e.g., Ref. [10]). For comparison, the black line also reports the value of $S_{L/2}$ for a fully random state of the form
|$\ket{\psi}=\frac{1}{\sqrt{\mathcal{N}_{L}}}\sum_{\{n_{j}\}}\operatorname{e}^{-i\varphi_{\{n_{j}\}}}\ket{\{n_{j}\}}\,,$ (24)
where |$\ket{\{n_{j}\}}$ are the simultaneous eigenstates of the operators $\hat{n}_{j}$ and $\varphi_{\{n_{j}\}}$ are random phases uniformly distributed in $[0,2\pi]$. This has been worked out some time ago by Page [102]. We find that $K$ closely follows the value predicted by Page, suggesting a thermal behavior of the half-system reduced density matrix in the limit $\gamma\to 0$, in agreement with previous results on systems obeying eigenstate thermalization [12, 13, 17, 11].
@@
Let us now comment on the behavior of $Q$. In fact, as clearly emerging from Fig. 6(c), it behaves quite differently for the two models. The double-logarithmic plot tells us that this is consistent with $Q\propto L^{y}$, a scaling we would have expected to get a behavior described by Eq. (1). Although the achievable sizes are too small for a large-$L$ extrapolation, we may obtain an estimate of the exponent $y$ by applying a linear fit to
@@
From the one side, for the SYK model we obtain a superlinear scaling $\overline{S}_{L/2}\sim L^{1.207}$ [cf Eq. (26b)]. Of course, a superlinear growth of the EE cannot be possible for arbitrarily large sizes and, in fact, it is due to finite-size effects. To corroborate this statement, in Fig. 7 we plot the asymptotic EE of the SYK model versus $L$, for different values of $\gamma$, and compare with the value predicted by Page [102] for a random state as in Eq. (24) (black line): after an initial superlinear transient, which can be better appreciated for small values of $\gamma$, all the curves approach a linear behavior that is below the Page value. This result shows that the fully chaotic nature of the SYK model [98, 99] (namely, all its eigenstates display volume-law entanglement [100, 101]) is so robust to survive the measurement process and to lead to a volume-law scaling of the entanglement with the size, independently of the measurement strength $\gamma$.
@@
From the other side, for the staggered $t$-$V$ model we find a very different behavior. Since $0.571-b\approx-$ 0.57 [cf Eq. (26a)], at some point the EE should start decreasing. This is likely to be ascribed to a finite-size effect: For larger sizes the fit with Eq. (20) might not work anymore. This is corroborated by the fact that, in this case, $\beta(L)$ increases with the system size [see Fig. (6)(a)] and does not saturate, so the correct form is $\overline{S}_{L/2}\sim L^{-0.57}/\gamma^{\beta(L)}$. This means that, for $\gamma<1$, the increase of $\beta$ might compensate the decrease of $L^{-0.57}$ and the area-law behavior might survive only for $\gamma>1$. Therefore, our results seem to suggest the presence of an entanglement transition from an area-law behavior, for $\gamma\gtrsim 1$, to a phase characterized by some kind of entanglement increase, for $\gamma\lesssim 1$. Unfortunately, our numerics does not allow us to make any precise statement on that.

### v2

@@ IV.1 Tight-binding chain with onsite dephasing
where $J$ denotes the hopping strength and $\hat{n}_{j}=\hat{c}^{\dagger}_{j}\hat{c}_{j}$ is the onsite fermion number operator. Here and in the other considered one-dimensional models, we adopt periodic boundary conditions by assuming $\hat{c}^{(\dagger)}_{L+1}\equiv\hat{c}^{(\dagger)}_{1}$. With reference to Eq. (8), we have $p_{j}=0$ and $q_{j}=1$. This system possesses a $U(1)$ symmetry, corresponding to the conservation of the total number of fermions, $\hat{N}=\sum_{j}\hat{n}_{j}$.
[…]
@@
Figure 1: The asymptotic averaged EE for the model in Eqs. (16). (a) Some examples with the behavior of $\overline{S}_{L/2}$ versus $L$ (circles) for different values of $\gamma$, together with the corresponding fits with Eq. (1) (continuous lines) where we fix the parameter $b$ as $b=1$. (b) The length scale $L_{0}=1/C$ —with $C$ obtained by fitting the curves in panel (a) with Eq. (1)— versus $\gamma$ in a double-logarithmic plot. The straight line results from the fit of the data for $\gamma<0.1$ and corresponds to a power law of the form $L_{0}\sim\gamma^{-0.88}$. We simulate the time evolution until $t_{f}=6\times 10^{6}$, with a step $\delta t=0.01$. Here and in the next figures, we work in units of $J=1$.
For the model in Eqs. (16), the authors of Ref. [46] showed the existence of an area-law phase for the asymptotic EE. More recently, the existence of a transition from area- to logarithm-law has been first claimed [54, 94] and then challenged. In fact, through the replica trick within a Keldysh path-integral formalism, it has been suggested that only the area-law phase exists, while the logarithm-law phase should be just a finite-size crossover, due to the exponential growth of a localization length with the inverse measurement strength [51]. For this reason, we fix $b=1$ in the fitting function (1), so that asymptotically for large system sizes one gets an area law.
Figure 1(a) displays numerical results for $\overline{S}_{L/2}$ versus $L$ (circles), for some values of the measurement strength $\gamma$, and the corresponding fit of Eq. (1) (continuous lines). We observe an accurate agreement between the two. This finding confirms the asymptotic area-law prediction, considering that the fit is meaningful because the maximum of the fitting range is comparable to the scale $L_{0}=1/C$ separating the short-range from the long-range behavior. This is evaluated using the value of $C$ extracted from the fit and is plotted versus $\gamma$ in a double-logarithmic plot [see Fig. 1(b)]. Notice that, for $\gamma<0.1$, the data tend to behave as a power law (straight line in the log plot). Applying a least-square linear fit we get $L_{0}\simeq\gamma^{-0.88}$. This result is not very far from the behavior $L_{0}\simeq 1/\gamma$ predicted in Ref. [46].
@@ IV.2 Kitaev chain with onsite dephasing
where $J$ is the nearest-neighbor coupling and $2h$ is a local chemical potential. This system has a $\mathbb{Z}_{2}$ symmetry, since the parity $\hat{P}=\prod_{j}\hat{n}_{j}$ of the fermion number is conserved (the number of particles $\hat{N}$ itself is not conserved, due to the presence of the pairing terms $\hat{c}_{j}^{\dagger}\hat{c}_{j+1}^{\dagger}$).
@@
where ${\cal N}_{t}$ is a normalization prefactor and ${Z}_{t}$ is an antisymmetric $L\times L$ matrix that can be written as ${Z}_{t}=-[{U}_{t}^{\dagger}]^{-1}\,{V}_{t}^{\dagger}$. The $U_{t}$ and $V_{t}$ can be cast as the submatrices of a Bogoliubov rotation allowing to construct the fermionic operators that annihilate the unraveled state (19) and obey linear differential equations [96, 77]. The interpretation as a Bogoliubov rotation is valid if ${U}_{t}$ and ${V}_{t}$ obey a unitarity condition, a constraint that can be restored (keeping ${Z}_{t}$ unchanged) by using a QR decomposition [77]. One can therefore restrict to study the dynamics of the two $L\times L$ matrices ${U}_{t}$ and ${V}_{t}$, keeping a polynomial scaling of the problem complexity and thus allowing the numerics to reach systems with a few hundreds of sites.
The monitored dynamics of the model in Eqs. (18) has been widely studied from a numerical point of view [55, 57, 58], always supporting a transition from a logarithm-law to an area-law regime, depending both on the measurement strength and on the parameter $h$ in the Hamiltonian (18a). This transition has been also analytically proved by exploiting an approach based on a nonlinear sigma model [50].
Figure 2: The $E$ for the model in Eqs. (18). (a) $\overline{S}_{L/4}$ versus $L$ (circles) for different values of $h$ and fixed $\gamma=1.5$, together with the corresponding fits with Eq. (1) (continuous lines). (b) The fitting parameter $b$ —obtained by fitting the curves in panel (a) with Eq. (1)— versus $h$. (c) The length scale $L_{0}$ [see Eq. (2)] versus $h$ obtained with the numerical fit. We evolve up to $t_{f}=60$ with a step $\delta t=0.05$ and take $N_{\rm r}=100$.
In Fig. 2(a) we show our numerical results for $\overline{S}_{L/2}$ versus $L$ (circles), for some values of $h$ and fixed $\gamma$, and the corresponding fit obtained using Eq. (1) (lines). Even in this case we observe a nice agreement between the numerical data and the fitting function, over all the considered range of sizes $16\leq L\leq 192$. In Fig. 2(c) we plot $L_{0}$ versus $h$ and see that it is always significantly smaller than the maximum of the range of sizes where we apply the fit $L_{\rm max}=256$ as shown in Fig. 2(a)], thus confirming the reliability of our fit. Looking in more detail at the fit exponent $b$ as a function of $h$ [Fig. 2(b)], we observe a monotonically increasing behavior, contrary to that for the tight-binding model reported in Fig. 1. We also see that $b$ now gets significantly smaller than one, for small values of $h$. So we can more confidently state that, in this case, there is a large-$h$ regime where the EE displays and area-law behavior and a small-$h$ regime where the EE is likely scaling with the system size in a sub-volume way. Unfortunately, it is difficult to mark a precise crossover point: in the region where $b$ is slightly smaller than one, the same issues occurring for the model of Sec. IV.1 emerge. In particular, with the available system sizes, it is impossible to distinguish between an area-law and a logarithm-law behavior. Curiously, the threshold value $h\approx 3$, conjectured to be the crossover point (for $\gamma=1.5$) on the basis of an alternative fit of the numerical data up to $L=256$ performed in Ref. [58], is compatible with the analysis reported in Fig. 2(b). However we stress that, despite the procedure in Ref. [58] was rather sensitive to finite-size effects, the one outlined here seems to us more appropriate and robust in this sense (see Appendix A for details on the numerical stability of the fits). Here we have only discussed the case $\gamma=1.5$, although we checked that analogous considerations apply for other values of the system-bath coupling, leading to the same qualitative conclusions (not shown).
@@ IV.3 Kitaev chain with long-range dissipators
[…]
Figure 3: The $E$ for the model in Eqs. (18a) and (LABEL:Z2_int_lr:eqn). (a) $\overline{S}_{L/2}$ versus $L$ (circles) for different values of $\alpha$ and fixed $\gamma=0.1$, $h=0.5$, together with the corresponding fits with Eq. (1) (continuous lines). (b) The fitting parameter $b$ —obtained by fitting the curves in panel (a) with Eq. (1)— versus $\alpha$. (c) The length scale $L_{0}$ [see Eq. (2)] versus $\alpha$ obtained with the numerical fit. We simulate the time evolution until $t_{f}=3\times 10^{3}$, with a step $\delta t=0.005$. The shaded areas locate the two crossover regions found in Ref. [77] for the same set of parameters used here.
All these regimes can be recognized by fitting the asymptotic averaged EE with Eq. (1). To show this fact, we concentrate on the case $\gamma=0.1$ and $h=0.5$. In Fig. 3(a) we show numerical data for different power-law exponents $\alpha$ (circles) and the corresponding fit (lines), that nicely reproduces all the curves. In Fig. 3(b) we show the parameter $b$ vs $\alpha$. The shaded areas locate the two crossover regions, respectively at $0.5\lesssim\alpha^{\star}_{1}\lesssim 1$ and at $\alpha^{\star}_{2}\approx 3.2$, discussed in Ref. [77]. We expect a volume-law behavior for $\alpha<\alpha^{\star}_{1}$ and, in fact, we find $b\approx 0$. In contrast, for $\alpha>\alpha^{\star}_{2}$, we expect an area-law regime and, in fact, we get $b>1$. Finally, in the intermediate region between area- and volume-law regimes where a subvolume increase in $L$ was observed, we consistently find an exponent $b\approx 0.8$, which is roughly constant in all the region. Regarding the last observation, we point out that in Ref. [77] the observed subvolume growth was faster than the logarithmic one. Moreover, no analytical function to describe the behavior of the entanglement with the system size was proposed there, while the fitting function Eq. (1) we suggest here well describes this behavior. Even for this model, qualitatively analogous considerations apply for other values of $\gamma$.
In Fig. 3(c) we plot the length scale $L_{0}$ given in Eq. (2) versus $\alpha$. In the range where we can evaluate it, we find it to be smaller than the maximum of the range where the fit is applied $L_{\rm max}=192$, see Fig. 3(a)], marking the reliability of the fit. For $\alpha\leq 1$ it diverges. This is consistent with the fact that, in this range of $\alpha$, the entanglement grows as a volume-law, thus the length scale $L_{0}$ separating the small-size range of linear increase from the large-size range of power-law increase in Eq. (1) entropy diverges. Notice that when $\alpha=1$ one has $b>0$ and $L_{0}$ is not divergent but finite, although much larger than $L_{\rm max}$. (We do not show it because it is of order $10^{10}$.) This value of $\alpha$ is probably the transition point between volume-law and power-law regimes; Anyway numerics confirms that here the steady-state entanglement entropy behaves linearly on the whole range where the fit is applied, being $L_{0}\gg L_{\rm max}$.
@@ V Nonintegrable models
In Sec. V.1 we describe how $\overline{S}_{\ell}$ versus $\gamma$ can be fitted by the function in Eq. (21) for the $t$-$V$ model with onsite dephasing, while in Sec. V.2 we do the same for the SYK model. In Sec. V.3, we discuss how the parameters $K,\,Q,\,\beta$ depend on the size $L$. This analysis shows that, in both cases, the dependence of $\overline{S}_{\ell}$ on $L$ is of the same form as in Eq. (1); this finding provides us with the rationale for fitting $\overline{S}_{\ell}$ versus $\gamma$ with a non intuitive function as Eq. (21). As a last stage to understand nonintegrable models, in Sec. V.4 we focus on their localization properties, by checking for the scaling of the IPR with the dimension of the Hilbert space. We find an anomalous delocalization behavior, which is apparently not related to the behavior of the EE.
@@ V.1 Staggered $t$-$V$ model with onsite dephasing
Figure 4: The $E$ for the model in Eqs. (22). Some examples of $\overline{S}_{L/2}$ versus $\gamma$ (circles), for various system sizes up to $L=20$ (see legend), together with the corresponding fits with the generalized Lorentzian function in Eq. (21) (continuous lines). We simulate the time evolution until $t_{f}=2\times 10^{3}$ with a step $\delta t=0.01$, while we set $W=V=1$.
Figure 4 displays our numerical results for the asymptotic averaged EE $\overline{S}_{L/2}$ versus the measurement strength $\gamma$ (circles) and the corresponding fit obtained with Eq. (21) (continuous lines). We can see that the latter performs well over a range of $\gamma\in[8\times 10^{-3},4]$ corresponding to more than two orders of magnitude.
@@ V.2 SYK model with onsite dephasing
The SYK Hamiltonian is a fermionic long-range interacting lattice model, being characterized by random four-particle interactions [99, 100]. Adding dissipation in the form of local dephasing, as in Eq. (16), the model can be written as
@@
Figure 5: The $E$ for the model in Eqs. (23). We show some examples of $\overline{S}_{L/2}$ versus $\gamma$ (circles), for various sizes up to $L=20$ (see legend), together with the corresponding fits with Eq. (21) (continuous lines). We simulate the time evolution until $t_{f}=3.4\times 10^{2}$, with a step $\delta t=0.01$, and fix $J=1$.
The SYK model is elusive to perturbative treatments at any energy scale, lying far outside the quasiparticle paradigm. In fact, it is known to be a paradigm for quantum chaos, displaying fast scrambling [104, 105], a nonzero entropy density at vanishing temperature [102], and a bipartite EE that scales linearly with $L$ for all the eigenstates (even for the ground state) [106, 107]. In the context of entanglement transitions, a related model of Brownian SYK chains subject to continuous monitoring has been considered in Ref. [88].
In Fig. 5 we show our numerical results for the averaged asymptotic EE versus the measurement strength (circles) and the fit obtained with Eq. (21) (continuous lines), displaying again a good agreement between the two, over the same range of $\gamma$ values as in the $t$-$V$ model.
@@ V.3 Discussion
The results showed in the two previous subsections cannot help too much in determining the asymptotic properties of the EE. However, some information can be deduced by looking at the fitting parameters. In Fig. 6, we show the behaviors of $\beta$ vs $L$ (a), $K$ vs $L$ (b), and of $\ln Q$ vs $\ln L$ (c), for both the $t$-$V$ (orange) and the SYK (green) models.
@@
we find $x=1.023\pm 0.008$ for the $t$-$V$ model, while $x=0.955\pm 0.039$ for the SYK model. Therefore $K$, corresponding to the value of the EE in the $\gamma\to 0$ limit, increases linearly with the system size $L$. For comparison, the black line also reports the value of $S_{L/2}$ for a fully random state of the form
[…]
@@
Let us now comment on the behavior of $Q$. In fact, as clearly emerging from Fig. 6(c), it behaves quite differently for the two models. The double-logarithmic plot tells us that this is consistent with $Q\propto L^{y}$, a scaling consistent with a behavior described by Eq. (1). Although the achievable sizes are too small for a large-$L$ extrapolation, we may obtain an estimate of the exponent $y$ by applying a linear fit to
@@
From the one side, for the SYK model we obtain a superlinear scaling $\overline{S}_{L/2}\sim L^{1.207}$ [cf Eq. (27b)]. Of course, a superlinear growth of the EE cannot be possible for arbitrarily large sizes and, in fact, it is due to finite-size effects. To corroborate this statement, in Fig. 7 we plot the asymptotic EE of the SYK model versus $L$, for different values of $\gamma$, and compare with the value predicted by Page [108] for a random state as in Eq. (25) (black line): after an initial superlinear transient, which can be better appreciated for small values of $\gamma$, all the curves approach a linear behavior that is below the Page value. This result shows that the fully chaotic nature of the SYK model [104, 105] (namely, all its eigenstates show an EE linear in the system size [106, 107]) is so robust to survive the measurement process and to lead to a linear increase of the steady-state entanglement with the size, independently of the measurement strength $\gamma$.
@@
From the other side, for the staggered $t$-$V$ model we find a very different behavior. Since $1-b\approx-0.57$ [cf Eq. (27a)], at some point the EE should start decreasing. This is likely to be ascribed to a finite-size effect: For larger sizes the fit with Eq. (21) might not work anymore. This is corroborated by the fact that, in this case, $\beta(L)$ increases with the system size [see Fig. (6)(a)] and does not saturate, so the correct form is $\overline{S}_{L/2}\sim L^{-0.57}/\gamma^{\beta(L)}$. This means that, for $\gamma<1$, the increase of $\beta$ might compensate the decrease of $L^{-0.57}$ and the area-law behavior might survive only for $\gamma>1$. Therefore, our results suggest the presence of an entanglement crossover from an area-law behavior, for $\gamma\gtrsim 1$, to a regime characterized by some kind of entanglement increase, for $\gamma\lesssim 1$. Unfortunately, our numerics does not allow us to make any precise statement on that.